Solve: x√121 = 11^x
I know that the answer is 2, but I don't understand how to reach the answer...

Answer by stanbon(75887)

Solve: x√121 = 11^x
I know that the answer is 2, but I don't understand how to reach the answer
----
x*11 = 11^x
----
x = 11^(x-1)
-----
The answer to that is not 2.
Cheers,
Stan H.